If the sound level in a room is increased from 50 dB to 60 dB, by what factor is the pressure amplitude increased?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. β base 1 = 50 dB, β base 2 = 60 dB ∴ I base 1 = 10^–7 W/m^2, I base 2 = 10^–6 W/m^2 (because β = 10 log base 10 (I/I base 0), where I base 0 = 10^–12 W/m^2) Again, I base 2/I base 1 = (p base 2/p base 1)^2 =(10^–6/10^–7) = 10 (where p = pressure amplitude). ∴ (p base 2 / p base 1) = √10 .